Circular No. 8789

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006sr–2006te Five supernovae have been reported from unfiltered CCD exposures:

2006ss by T. Boles (cf. IAUC 8781), 2006st by D. Winslow and W. Li (LOSS/KAIT; cf. IAUC 8782), 2006su by G. Duszanowicz (cf. IAUC 8771), 2006td by W. Kloehr (cf. IAUC 8553), and 2006te by O. Trondal, P. Luckas, and M. Schwartz (0.35-m telescope, Oslo; cf. IAUC 8733). Also, J. Frieman and the Sloan Digital Sky Survey II collaboration have reported eight faint new type-Ia supernovae (designated 2006sv–2006tc; magnitude range g = 21.9–22.7) found in November, with details given on CBET 786.

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ss Dec. 17.21 14^{h}20^{m}27.^{s}46 +35^{o}11′42′′.7 17.6 11′′.9 E, 22′′.7 N
2006st Dec. 18.57 11 58 40.97 +27 27 37.6 18.3 63′′ W, 29′′ N
2006su Dec. 18.15 12 46 56.88 +36 32 01.8 17.7 8′′.5 E, 10′′.5 N
2006td Dec. 24.99 1 58 15.76 +36 20 57.7 15.1 4′′.5 N
2006te Dec. 28.91 8 11 42.99 +41 33 16.8 15.8 5′′.5 W, 1′′.7 S

Additional magnitudes for 2006ss in NGC 5579: 1988 Mar. 16, [21.0 (Digitized Sky Survey, blue); 1992 May 1, [20.5 (DSS, red); 2006 Mar. 6 and June 30, [19.5 (Boles); Dec. 18.111 UT, 17.6 (Boles). SN 2006ss is a type-IIb supernova, ∼ 19 days past explosion on Dec. 21 (spectroscopic details on CBET 781). Additional KAIT magnitudes for 2006st in NGC 4017: May 30.23, [19.0; Dec. 19.53, 18.4. SN 2006st is a type-II supernova (details on CBET 778). Additional magnitudes for 2006su in IC 3772: 2005 Apr. 11.92, [18.0; 2006 Dec. 20.166, 17.4. SN 2006su is a type-Ia supernova about a month past maximum (details on CBET 784). Additional magnitudes for 2006td: 2002 Sept. 10.97, [19.5 (O. Trondal, Oslo, Norway; composite ‘RGB’ frame); 2006 Dec. 25.89, 15.0 (G. Masi, Ceccano, Italy; source of the precise position tabulated above; unfiltered CCD); 25.914, 15.6 (P. Corelli, Pagnacco, Italy; unfiltered CCD). Nothing is visible at the position of 2006td on a Palomar Sky Survey image (limiting magnitude I ∼ 19). SN 2006td is a type-Ia supernova (spectroscopic details on CBET 790). Additional magnitudes for 2006te: 1989 Nov. 30, [21.0 (Digitized Sky Survey, red); 2006 Jan. 28.97, [18.0 (Trondal); Dec. 30.88, 15.8 (E. Konstantinos, 0.30-m telescope + CCD). Also, 2006sr (cf. IAUC 8784) is a type-Ia supernova discovered a few days prior to maximum (spectroscopic details on CBET 779).

2006 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green